OpenStudy (anonymous):

will give medal! what is x? -6.8+2.6+x=9.7

OpenStudy (anonymous):

can you help me with this one @ayoubEpst ?

OpenStudy (anonymous):

x=9.7+6.8-2.6

OpenStudy (anonymous):

when you move it to the other side change the sign

OpenStudy (anonymous):

13.9

OpenStudy (anonymous):

ohh okay thank you can you help me with one more

OpenStudy (anonymous):

13.9

OpenStudy (anonymous):

so x= 13.9

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }(3x-16)+\frac{ 1 }{ 4 }x= 5 \frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

@ayoubEpst this is the last one

OpenStudy (anonymous):

x=5/3 +4

OpenStudy (anonymous):

x=17/3

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

if you want i can explain how i find it ?